OpenStudy (anonymous):

1/(x+a) = 2/(x-1) - 1/(x-a). Solve for x? (where a,b are constants)

myininaya (myininaya):

i seen this problem the other day let me see if i can find that question

OpenStudy (zarkon):

"where a,b are constants"...there are no b's in your equation

OpenStudy (anonymous):

yes b is missing

OpenStudy (anonymous):

i think it says that just if there happen to be a b. but yes there is no b so just ignore that part. *given that a is constant*

OpenStudy (anonymous):

can u plz rectify ur question

OpenStudy (anonymous):

x=a

OpenStudy (anonymous):

myininaya are you sure it couldnt be simpler?

OpenStudy (anonymous):

yeah i remember this one. was happy for you to do it!

OpenStudy (anonymous):

x=a^2

OpenStudy (anonymous):

x=a^2

OpenStudy (anonymous):

start with \[\frac{1}{x+a}=\frac{2}{x-1}-\frac{1}{x-a}\] multiply all this nonsense on both sides by \[(x+a)(x-1)(x-a)\] do a raft of algebra and then get \[x=a^2\]

OpenStudy (anonymous):

yeah right u r.i missed it.

OpenStudy (anonymous):

if you need all the annoying algebra steps, click on "show steps" http://www.wolframalpha.com/input/?i=solve+for+x+1%2F%28x%2Ba%29+%3D+2%2F%28x-1%29+-+1%2F%28x-a%29.

OpenStudy (anonymous):

by simplifying lcd=(x-1)(x+a)(x-a) (x-1)(x-a)-2[(x+a)(x-a)]+(x-1)(x-a)=0 2x^2+a-x-a-2x^2+a^2-ax-x+a=0 x=a^2